# Adaptive Grid-Based Thompson Sampling for Efficient Trajectory Discovery

### v1

@@ Abstract
Bayesian optimization (BO) is a powerful framework for estimating parameters of computationally expensive simulation models, particularly in settings where the likelihood is intractable and evaluations are costly. In stochastic models every simulation is run with a specific parameter set and an implicit or explicit random seed, where each parameter set and random seed combination generates an individual realization, or trajectory, sampled from an underlying random process. Existing BO approaches typically rely on summary statistics over the realizations, such as means, medians, or quantiles, potentially limiting their effectiveness when trajectory-level information is desired. We propose a trajectory-oriented Bayesian optimization method that incorporates a Gaussian process (GP) surrogate using both input parameters and random seeds as inputs, enabling direct inference at the trajectory level. Using a common random number (CRN) approach, we define a surrogate-based likelihood over trajectories and introduce an adaptive Thompson Sampling algorithm that refines a fixed-size input grid through likelihood-based filtering and Metropolis-Hastings-based densification. This approach concentrates computation on statistically promising regions of the input space while balancing exploration and exploitation. We apply the method to stochastic epidemic models, a simple compartmental and a more computationally demanding agent-based model, demonstrating improved sampling efficiency and faster identification of data-consistent trajectories relative to parameter-only inference.
@@ 1 Introduction
Epidemiological simulation models are increasingly recognized as indispensable tools for supporting public health decisions during crises (e.g., Ozik2021 ; eubank2004modelling ). Their use dates back to at least the early twentieth century, when researchers began modeling how infectious diseases—including influenza and, more recently, COVID-19—propagate through contact networks. The most common types of these models, such as compartmental models and agent-based models (ABMs), each have a set of parameters governing disease dynamics. Some models are deterministic, producing identical outputs for a given set of parameter inputs. Others are stochastic, generating outputs based on both model parameters and random seeds, producing specific stochastic realizations, or trajectories, of random processes within the models. Estimating model parameters typically involves fitting simulated trajectories to observed data (chatzilena2019contemporary, ; toni2009approximate, ; pooley2015posterior, ). In this paper we investigate a novel Bayesian optimization (BO) approach that seeks to align simulation outputs with empirical data at the trajectory level. This contrasts with more common approaches that rely on summary statistics over stochastic realizations.
@@
Classical calibration approaches rely on maximum likelihood estimation or Bayesian inference, often facilitated by surrogate models such as Gaussian processes (GPs) to emulate the expensive simulator and reduce computational cost. BO builds on this idea by embedding the surrogate in a sequential decision framework to prioritize promising regions of the parameter space. Other methods such as Ensemble Kalman Filtering (EnKF) (evensen2003ensemble, ; iglesias2013ensemble, ), the Partially Observed Markov Process (POMP) framework (king2016statistical, ), Approximate Bayesian Computation (ABC) (beaumont2002approximate, ), and simulation-based inference (SBI) (cranmerFrontierSimulationbasedInference2020, ; dyerBlackboxBayesianInference2024a, ) are also used extensively in calibrating simulation models. However, calibrating stochastic simulation models remains a fundamentally ill-posed problem (baker2022analyzing, ). Unlike deterministic models, where each input maps to a fixed output, stochastic simulators produce outputs that are random draws from an unknown distribution. The same input parameter $\mathbf{x}$ can yield divergent outcomes across simulation replicates, making it ambiguous what it means for parameters to match observed data. Nonetheless, in common practice, calibration involves aligning expected values or other distributional summaries of stochastic simulation trajectories with data.
@@
For the purpose of illustration, let us consider the use of ABMs for modeling infectious diseases. In an ABM, each individual is represented by an autonomous agent with predefined attributes and behaviors (eubank2004modelling, ). Agents interact over network structures that capture heterogeneous mixing patterns, and a single ABM run produces one possible mixing and transmission trajectory (ferguson2020report, ; chang2021modelling, ). Stochasticity arises from probabilistic transmission events but also from structural variations such as dynamic contact patterns or initial conditions. This results in different outcomes even under identical model parameters. When stochasticity reflects intrinsic structural variation, knowing only the parameter values that match observations ignores specific mixing patterns underlying transmission, which may be essential for planning targeted interventions. From a statistical perspective, such aggregation over randomness induces a loss of identifiability. Hence, we argue that a principled solution should investigate the benefits of estimating both the parameters $\mathbf{x}$ and specific stochastic realizations, denoted by random seeds $r$, which jointly produce trajectories consistent with observed data. Recovering this tuple $(\mathbf{x},r)$ then enables trajectory-level inference. Individual trajectories also enable the use of trajectory continuations when modeling epidemics over time with updating surveillance data, such as with sequential data assimilation approaches (spannaus2022inferring, ; fadikar2024towards, ).
@@
Trajectory-oriented optimization (TOO) is a surrogate-based framework for calibrating stochastic simulations by jointly estimating the tuple $(\mathbf{x},r)$ with the use of a common random number GP (CRNGP) (fadikar2023TOO, ). The CRNGP enables the treatment of the simulation as a deterministic function over the augmented input space $(\mathbf{x},r)$, enabling inference at the trajectory level rather than through distributional summaries. Building on the TOO framework, this paper focuses on improving its computational efficiency to enable faster identification of simulated trajectories that are consistent with the observed data. We extend the Thompson Sampling (TS) based BO algorithm by adaptively generating the parameter-seed space for faster trajectory exploration. We illustrate this approach using a Susceptible-Exposed-Infectious-Recovered (SEIR) model of COVID-19 transmission in Chicago, highlighting the efficiency gains over existing trajectory-finding methods. We also discuss ways in which high-performance computing (HPC) workflows can extend the applicability of such complex algorithms to more computationally demanding simulations.
@@ 2.1 Gaussian Process Surrogates
GPs (Rasmussen2005, ; gramacy2020surrogates, ) are nonparametric regression models used to infer unknown functions from observed data. GPs have been used extensively as surrogates (Sacks1989a, ; higdon2004combining, ) for expensive deterministic computer simulations, for tasks such as sensitivity analysis (saltelli2008global, ), optimization (pearce2022bayesian, ), calibration (Kennedy2001, ), and uncertainty quantification (binois2015quantifying, ). Formally, a GP defines a distribution over functions $f:\mathbb{R}^{d}\to\mathbb{R}$, such that for any finite set of input locations $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}\subset\mathbb{R}^{d}$, the corresponding function evaluations $\{f(\mathbf{x}_{1}),\dots,f(\mathbf{x}_{N})\}$ follow a multivariate normal distribution. As a prior over functions, a GP is fully characterized by a mean function $\mu(\cdot)$ and a positive semi-definite covariance kernel $k(\cdot,\cdot)$, which encode assumptions about smoothness, stationarity, and correlation structure of the underlying function. Let $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ denote a training dataset consisting of $N$ pairs of inputs $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and corresponding scalar outputs $y_{i}\in\mathbb{R}$, where we assume $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ and $\epsilon_{i}\sim\mathcal{N}(0,\tau^{2})$ represents independent observation noise. Then, under a GP prior, the vector of outputs $\mathbf{Y}=(y_{1},\dots,y_{N})^{\top}$ is jointly Gaussian:
@@ 2.2 Tracking the Input Dependent Variability
While standard GP is a stationary process, and assumes homoskedastic noise, stochastic simulation models exhibit input-dependent variability that must be explicitly modeled. When replicate observations (multiple runs at the same design point) are available, the standard GP can be extended by aggregating replicate data and incorporating heteroskedastic noise into the covariance structure. One such extension is stochastic kriging (ankenmanstochastickriging, ), which allows for input-dependent noise modeling by incorporating known or estimated variance at each input. This approach explicitly accounts for variability in simulation outputs at different input locations. An advancement over stochastic kriging is the heteroskedastic Gaussian Process (hetGP) model (e.g., binois2018heteroscedastic, ), which jointly models both the mean and variance processes within a hierarchical framework. Unlike stochastic kriging, hetGP treats the variance as a latent function and estimates it from data, allowing for more flexibility in capturing input-dependent noise. However, it relies on the assumption that replicate variability can be adequately captured by a Gaussian distribution.
An alternative approach, the quantile Gaussian Process (QGP) (Plumlee2014, ; Fadikar2018, ), models conditional quantiles of the output distribution rather than its mean or variance. QGP makes no distributional assumptions and provides a richer characterization of uncertainty, particularly when simulation outputs exhibit asymmetry or heavy tails. These methods differ in their assumptions and objectives: stochastic kriging incorporates known variance estimates, hetGP infers them as latent processes under Gaussian assumptions, and QGP directly targets the shape of the output distribution.
@@ 2.3 Motivation for Trajectory-oriented Approaches
Figure 1 shows the calibrated deaths and hospitalizations trajectories from the CityCOVID model reported in Ozik2021 . For each input parameter, 30 stochastic replicates are shown (grey), together with their mean trajectory (blue). The figure illustrates that while mean trajectories can align well with the observed data (black dots), the individual realizations underlying these means can vary substantially, with many diverging away from the empirical outcomes. This variability highlights the influence of replicates in stochastic simulation models, where distinct realizations under identical parameters may produce very different epidemic trajectories. Consequently, calibration approaches that rely solely on matching mean behavior can conceal the presence of poor-quality trajectories and may fail to capture the specific realizations that are consistent with observed dynamics. This example serves as the primary motivation behind our goal of developing trajectory-oriented calibration approaches, which explicitly target the joint estimation of parameters and stochastic realizations, with the objective of identifying individual trajectories that align closely with the data. This also motivates us to look at the effectiveness of the calibration algorithm and the quality of the discovered trajectories through the lens of errors of the individual trajectories by the means of RMSE (root mean-squared error). More discussion follows in Section 4.3.
@@
Figure 1: Deaths and hospitalizations from CityCOVID under 10 parameterizations identified in Ozik2021 . For each parameterization, 30 replicates (grey) and their mean (blue) are shown against observed data (black dots).
@@ 2.4 Common Random Number Gaussian Process (CRNGP)
To enable trajectory-level inference over parameter-seed pairs, we adopt the Common Random Number Gaussian Process (CRNGP), a surrogate model designed for stochastic simulations (fadikar2023TOO, ). The CRNGP models each simulation output using an augmented input pair $(\mathbf{x},r)$, where $\mathbf{x}$ is the input parameter and $r$ is a replicate identifier (e.g., we use the random seed for initializing a simulation’s random stream as the replicate identifier in this work), treated as a categorical variable (chen2012effects, ; pearce2019bayesian, ; pearce2022bayesian, ). Formally, we define $f:\mathbb{R}^{d}\times\mathcal{S}\to\mathbb{R}$, where $\mathcal{S}$ denotes a countable set of replicates. The simulator is assumed deterministic over the joint input $(\mathbf{x},r)$, and a standard GP prior is placed on $f$ with a separable kernel structure:
@@
The resulting kernel $k$ is positive definite and the inter-replicate similarity is controlled via $\rho$. A nugget term $\tau^{2}$ is added to the diagonal to form the full covariance $\mathbf{K}_{N}+\tau^{2}\mathbf{I}$. The kriging equations from Equation 2 remain applicable, and predictions can be made at any $(\mathbf{x},r)$. Averaging over seeds corresponds to predicting at a new, unobserved replicate (pearce2022bayesian, ).

### v2

# Staying on Track: Efficient Trajectory Discovery with Adaptive Batch Sampling

@@ Abstract
Bayesian optimization (BO) is a powerful framework for estimating parameters of expensive simulation models, particularly in settings where the likelihood is intractable and evaluations are costly. In stochastic models every simulation is run with a specific parameter set and an implicit or explicit random seed, where each parameter set and random seed combination generates an individual realization, or trajectory, sampled from an underlying random process. Existing BO approaches typically rely on summary statistics over the realizations, such as means, medians, or quantiles, potentially limiting their effectiveness when trajectory-level information is desired. We propose a trajectory-oriented BO method that incorporates a Gaussian process surrogate using both input parameters and random seeds as inputs, enabling direct inference at the trajectory level. Using a common random number approach, we define a surrogate-based likelihood over trajectories and introduce an adaptive Thompson Sampling algorithm that refines a fixed-size input grid through likelihood-based filtering and Metropolis-Hastings-based densification. This approach concentrates computation on statistically promising regions of the input space while balancing exploration and exploitation. We apply the method to stochastic epidemic models, a simple compartmental and a more computationally demanding agent-based model, demonstrating improved sampling efficiency and faster identification of data-consistent trajectories relative to parameter-only inference.
@@ 1 Introduction
Epidemiological simulation models are increasingly recognized as indispensable tools for supporting public health decisions during crises (e.g., Ozik et al. (2021); Eubank et al. (2004)). Their use dates back to at least the early twentieth century, when researchers began modeling how infectious diseases including influenza and, more recently, COVID-19, propagate through contact networks (Kermack and McKendrick, 1927). The most common types of these models, such as compartmental models and agent-based models (ABMs), each have a set of parameters governing disease dynamics. Some models are deterministic, producing identical outputs for a given set of parameter inputs. Others are stochastic, generating outputs based on both model parameters and random seeds, producing specific stochastic realizations, or trajectories, of random processes within the models. Estimating model parameters typically involves fitting simulated trajectories to observed data (Chatzilena et al., 2019; Toni et al., 2009; Pooley et al., 2015). In this paper we investigate a novel Bayesian optimization (BO) approach that seeks to align simulation outputs with empirical data at the trajectory level. This contrasts with more common approaches that rely on summary statistics over stochastic realizations.
@@
Classical calibration approaches rely on maximum likelihood estimation or Bayesian inference, often facilitated by surrogate models such as Gaussian processes (GPs) to emulate the expensive simulator and reduce computational cost. BO builds on this idea by embedding the surrogate in a sequential decision framework to prioritize promising regions of the parameter space. Other methods such as Ensemble Kalman Filtering (EnKF) (Evensen, 2003; Iglesias et al., 2013), the Partially Observed Markov Process (POMP) framework (King et al., 2016), Approximate Bayesian Computation (ABC) (Beaumont et al., 2002), and simulation-based inference (SBI) (Cranmer et al., 2020; Dyer et al., 2024) are also used extensively in calibrating simulation models. However, calibrating stochastic simulation models remains a fundamentally ill-posed problem (Baker et al., 2022). Unlike deterministic models, where each input maps to a fixed output, stochastic simulators produce outputs that are random draws from an unknown distribution. The same input parameter $\mathbf{x}$ can yield divergent outcomes across simulation replicates, making it ambiguous what it means for parameters to match observed data. Nonetheless, in common practice, calibration involves aligning expected values or other distributional summaries of stochastic simulation trajectories with data.
@@
For the purpose of illustration, let us consider the use of ABMs for modeling infectious diseases. In an ABM, each individual is represented by an autonomous agent with predefined attributes and behaviors (Eubank et al., 2004). Agents interact over network structures that capture heterogeneous mixing patterns, and a single ABM run produces one possible mixing and transmission trajectory (Ferguson et al., 2020; Chang et al., 2021). Stochasticity arises from probabilistic transmission events but also from structural variations such as dynamic contact patterns or initial conditions. This results in different outcomes even under identical model parameters. When stochasticity reflects intrinsic structural variation, knowing only the parameter values that match observations ignores specific mixing patterns underlying transmission, which may be essential for planning targeted interventions. From a statistical perspective, such aggregation over randomness induces a loss of identifiability. Hence, we argue that a principled solution should investigate the benefits of estimating both the parameters $\mathbf{x}$ and specific stochastic realizations, denoted by random seeds $r$, which jointly produce trajectories consistent with observed data. Recovering this tuple $(\mathbf{x},r)$ then enables trajectory-level inference. Essentially, rather than characterizing stochasticity as an inconvenient model behavior to be averaged out or minimized in expectation, trajectory-level inference characterizes it as a feature that can aid in explanation of observed dynamics in conjunction with model parameters. This more deliberate treatment may also be interpreted as constructing a set of candidate realizations of a stochastic process that explain observed data instead of being interpreted as the estimated expected outcome under infinite realizations. Individual trajectories also enable the use of trajectory continuations when modeling epidemics over time with updating surveillance data, such as with sequential data assimilation approaches (Spannaus et al., 2022; Fadikar et al., 2024).
@@
Trajectory-oriented optimization (TOO) is a surrogate-based framework for calibrating stochastic simulations by jointly estimating the tuple $(\mathbf{x},r)$ with the use of a common random number GP (CRNGP) (Fadikar et al., 2023). The CRNGP enables the treatment of the simulation as a deterministic function over the augmented input space $(\mathbf{x},r)$, enabling inference at the trajectory level rather than through distributional summaries. Building on the TOO framework, this paper focuses on improving its computational efficiency to enable faster identification of simulated trajectories that are consistent with the observed data. We extend the Thompson Sampling (TS) based BO algorithm by adaptively generating the parameter-seed space for faster trajectory exploration. We illustrate this approach using a Susceptible-Exposed-Infectious-Recovered (SEIR) model of COVID-19 transmission in Chicago, highlighting the efficiency gains over existing trajectory-finding methods. We also discuss ways in which high-performance computing (HPC) workflows can extend the applicability of such complex algorithms to more computationally demanding simulations.
@@ 2.1 Gaussian Process Surrogates
GPs (Rasmussen and Williams, 2005; Gramacy, 2020) are nonparametric regression models used to infer unknown functions from observed data. GPs have been used extensively as surrogates (Sacks et al., 1989; Higdon et al., 2004) for expensive deterministic computer simulations, for tasks such as sensitivity analysis (Saltelli et al., 2008), optimization (Pearce et al., 2022), calibration (Kennedy and O’Hagan, 2001), and uncertainty quantification (Binois et al., 2015). Formally, a GP defines a distribution over functions $f:\mathbb{R}^{d}\to\mathbb{R}$, such that for any finite set of input locations $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}\subset\mathbb{R}^{d}$, the corresponding function evaluations $\{f(\mathbf{x}_{1}),\dots,f(\mathbf{x}_{N})\}$ follow a multivariate normal distribution. As a prior over functions, a GP is fully characterized by a mean function $\mu(\cdot)$ and a positive semi-definite covariance kernel $k(\cdot,\cdot)$, which encode assumptions about smoothness, stationarity, and correlation structure of the underlying function. Let $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ denote a training dataset consisting of $N$ pairs of inputs $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and corresponding scalar outputs $y_{i}\in\mathbb{R}$, where we assume $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ and $\epsilon_{i}\sim\mathcal{N}(0,\tau^{2})$ represents independent observation noise. Then, under a GP prior, the vector of outputs $\mathbf{Y}=(y_{1},\dots,y_{N})^{\top}$ is jointly Gaussian:
@@ 2.2 Tracking the Input Dependent Variability
While standard GP is a stationary process, and assumes homoskedastic noise, stochastic simulation models exhibit input-dependent variability that must be explicitly modeled. When replicate observations (multiple runs at the same design point) are available, the standard GP can be extended by aggregating replicate data and incorporating heteroskedastic noise into the covariance structure. One such extension is stochastic kriging (Ankenman et al., 2010), which allows for input-dependent noise modeling by incorporating known or estimated variance at each input. This approach explicitly accounts for variability in simulation outputs at different input locations. An advancement over stochastic kriging is the heteroskedastic Gaussian Process (hetGP) model (e.g., Binois et al., 2018), which jointly models both the mean and variance processes within a hierarchical framework. Unlike stochastic kriging, hetGP treats the variance as a latent function and estimates it from data, allowing for more flexibility in capturing input-dependent noise. However, it relies on the assumption that replicate variability can be adequately captured by a Gaussian distribution.
An alternative approach, the quantile Gaussian Process (QGP) (Plumlee and Tuo, 2014; Fadikar et al., 2018), models conditional quantiles of the output distribution rather than its mean or variance. QGP makes no distributional assumptions and provides a richer characterization of uncertainty, particularly when simulation outputs exhibit asymmetry or heavy tails. These methods differ in their assumptions and objectives: stochastic kriging incorporates known variance estimates, hetGP infers them as latent processes under Gaussian assumptions, and QGP directly targets the shape of the output distribution.
@@ 2.3 Motivation for Trajectory-oriented Approaches
Outcome variation for stochastic simulation models is driven by both changes to the model parameters as well as stochasticity. To explicate these two ideas, we first consider a simple example of an agent-based SIR model on a discrete two-dimensional grid with dimensions $0,50]^{2}$. Agents move randomly on the grid, and Infected agents may infect Susceptible neighboring agents with probability $\beta$ and remain infected for $\gamma$ time steps. In some sense, this is a similar model to the well-known SIR system of ordinary differential equations (which is one form of a wider class of compartmental models). If we vary the per-contact transmission probability $\beta$ (which serves a similar role to the infection rate parameter in the compartmental approach), we expect a higher level of disease spread in the population. However, the agent-based simulation has an important caveat: it does not assume a “well-mixed” population like the compartmental model, which leads to path-dependence in the model outcomes (Hammond, 2015): an agent can only become Infected if one of its neighbors was already Infected at the previous time step, creating a chain of infection events. This also means that the $\beta$ parameter interacts nonlinearly with the agents’ movements. Therefore, there are likely some specific infection parameter values that create just enough infections to lead to disease outbreaks for some random seeds, but die off in other cases.
To demonstrate this, we conduct the following experiment: we instantiate a population of $N=2000$ agents with N1N-1 non-infected (Susceptible) agents at unique locations on the grid and an index Infected agent near the top right corner. We vary the per-contact transmission probability as well as the random seed, the results of which are shown in Figure 1. The top left-hand panel demonstrates the effect of increasing the $\beta$ parameter and how it leads to increased population-level infection, where the replicate is fixed across the $\beta$ values. We also note that if the model at hand was deterministic, this would be the only way to induce changes to the model outcomes. The top right-hand panel demonstrates the specific effect of stochasticity: for a fixed $\beta$ =0.068=0.068 (corresponding to the $\beta$ parameter of the red line in the left panel of Figure 1a), cumulative infection also varies widely across random seeds. This highlights that in some simulations the epidemic fizzles out before infecting enough individuals to spread widely, and in some, takes off due to the idiosyncratic mixing that occurs in those specific replicates. These trajectories were selected so that each set induced a similar mean number of cumulative infections (shown as black dashed lines in both Figure 1a panels). Additionally, we see in the bottom panels (Figure 1b), where we track the correlations across random seeds, that the structure of the noise changes both across replicates and $\beta$ values. This supports the intuition that under certain conditions, information may be shared across stochastic replicates, in a similar fashion to how model parameters are learned through experimentation. This is relevant for active learning or BO workflows, where model runs are conducted sequentially, and whether a particular random seed was “good” or “bad” (the definition of which depends on the experiment’s goal) would be valuable information for subsequent sampling of the parameter (and replicate) space.
Figure 1: Variation in simulation outcomes due to changes in model parameters as well as stochasticity. Figure (a) shows cumulative epidemic curves for a simple SIR model of $2000$ agents on a grid, where the left-hand panel varies the $\beta$ parameter and fixes the random seed, while the right-hand panel fixes $\beta$ and varies the seed. Black dashed lines indicate the mean across trajectories. The red line in the left-hand panel denotes the trajectory for $\beta=0.068$, which is the fixed $\beta$ in right-hand panel. Figure (b) shows the squared correlation of cumulative infection curves across random seeds for two different values of $\beta$: when $\beta=0.068$ there is a high amount of variance between replicates, leading to widely different simulation outcomes, but when $\beta$ is much higher ($0.25$) the epidemic curves resemble each other much more closely across simulations. Each (x,y) point shows the squared correlation between two random seeds ordered from 1 to 50.
@@
Figure 2 extends this intuition beyond a simple pedagogical model and applies it to a real-world model that informed high-stakes decision-making: it shows the calibrated deaths and hospitalizations trajectories from the CityCOVID model reported in Ozik et al. (2021). For each input parameter, 30 stochastic replicates are shown (grey), together with their mean trajectory (blue). The figure illustrates that while mean trajectories can align well with the observed data (black dots), the individual realizations underlying these means can vary substantially, with many diverging away from the empirical outcomes. This variability highlights the influence of replicates in stochastic simulation models, where distinct realizations under identical parameters may produce very different epidemic trajectories. Consequently, calibration approaches that rely solely on matching mean behavior can conceal the presence of poor-quality trajectories and may fail to capture the specific realizations that are consistent with observed dynamics. This example serves as the primary motivation behind our goal of developing trajectory-oriented calibration approaches, which explicitly target the joint estimation of parameters and stochastic realizations, with the objective of identifying individual trajectories that align closely with the data. This also motivates us to look at the effectiveness of calibration algorithms and the quality of the discovered trajectories through the lens of errors of the individual trajectories by the means of RMSE (root mean-squared error). More discussion follows in Section 4.3.
@@
Figure 2: Cumulative deaths and incident hospitalizations from CityCOVID under 10 parameterizations identified in Ozik et al. (2021). For each parameterization, 30 replicates (grey) and their mean (blue) are shown against observed data (black dots).
@@ 2.4 Common Random Number Gaussian Process (CRNGP)
To enable trajectory-level inference over parameter-seed pairs, we adopt the Common Random Number Gaussian Process (CRNGP), a surrogate model designed for stochastic simulations (Fadikar et al., 2023). The CRNGP models each simulation output using an augmented input pair $(\mathbf{x},r)$, where $\mathbf{x}$ is the input parameter and $r$ is a replicate identifier (e.g., we use the random seed for initializing a simulation’s random stream as the replicate identifier in this work), treated as a categorical variable (Chen et al., 2012; Pearce et al., 2019, 2022). Formally, we define $f:\mathbb{R}^{d}\times\mathcal{S}\to\mathbb{R}$, where $\mathcal{S}$ denotes a countable set of replicates. The simulator is assumed deterministic over the joint input $(\mathbf{x},r)$, and a standard GP prior is placed on $f$ with a separable kernel structure:
@@
The resulting kernel $k$ is positive definite and the inter-replicate similarity is controlled via $\rho$. A nugget term $\tau^{2}$ is added to the diagonal to form the full covariance $\mathbf{K}_{N}+\tau^{2}\mathbf{I}$. The kriging equations from Equation 2 remain applicable, and predictions can be made at any $(\mathbf{x},r)$. Averaging over seeds corresponds to predicting at a new, unobserved replicate (Pearce et al., 2022).